# Open X-Embodiment: Robotic Learning Datasets and RT-X Models

### 18 Dec 2023 v4

@@ Abstract
Large, high-capacity models trained on diverse datasets have shown remarkable successes on efficiently tackling downstream applications. In domains from NLP to Computer Vision, this has led to a consolidation of pretrained models, with general pretrained backbones serving as a starting point for many applications. Can such a consolidation happen in robotics? Conventionally, robotic learning methods train a separate model for every application, every robot, and even every environment. Can we instead train “generalist” X-robot policy that can be adapted efficiently to new robots, tasks, and environments? In this paper, we provide datasets in standardized data formats and models to make it possible to explore this possibility in the context of robotic manipulation, alongside experimental results that provide an example of effective X-robot policies. We assemble a dataset from different robots collected through a collaboration between institutions, demonstrating skills ( tasks). We show that a high-capacity model trained on this data, which we call RT-X, exhibits positive transfer and improves the capabilities of multiple robots by leveraging experience from other platforms. More details can be found on the project website robotics-transformer-x.github.io.
@@ I Introduction
A central lesson from recent advances in machine learning and artificial intelligence is that large-scale learning from broad and diverse datasets can enable capable AI systems by providing for general-purpose pretrained models. In fact, large-scale general-purpose models typically trained on large and diverse datasets can often outperform their narrowly targeted counterparts trained on smaller but more task-specific data. For instance, open-vocabulary image classifiers (e.g., CLIP [1]) trained on large datasets scraped from the web tend to outperform fixed-vocabulary models trained on more limited datasets, and large language models [2, 3] trained on massive text corpora tend to outperform systems that are only trained on narrow task-specific datasets. Increasingly, the most effective way to tackle a given narrow task (e.g., in vision or NLP) is to adapt a general-purpose model. However, these lessons are difficult to apply in robotics: any single robotic domain might be too narrow, and while computer vision and NLP can leverage large datasets sourced from the web, comparably large and broad datasets for robotic interaction are hard to come by. Even the largest data collection efforts still end up with datasets that are a fraction of the size and diversity of benchmark datasets in vision (5-18M) [4, 5] and NLP (1.5B-4.5B) [6, 7]. Perhaps more importantly, such datasets are often still narrow along some axes of variation, either focusing on a single environment, a single set of objects, or a narrow range of tasks. How can we overcome these challenges in robotics and move the field of robotic learning toward the kind of large data regime that has been so successful in other domains?
@@
Following this rationale, our work has two primary goals: (1) Demonstrate that policies trained on data from many different robots and environments enjoy the benefits of positive transfer, attaining better performance than policies trained only on data from each evaluation setup. (2) Provide datasets, data formats and models for the robotics community to enable future research on X-embodiment models.
@@ V-A In-distribution performance across different embodiments
To assess the ability of our RT-X model variants to learn from X-embodiment data, we evaluate their performance on in-distribution tasks. We split our evaluation into two types of use cases: evaluation on domains that only have small-scale datasets (Fig. 2), where we would expect transfer from larger datasets to significantly improve performance, and evaluation on domains that have large-scale datasets (Table I), where we expect further improvement to be more challenging. Note that we use the same robotics data training mixture (defined in Sec. IV-C) for all the evaluations presented in this section. For small-scale dataset experiments, we consider Kitchen Manipulation [128], Cable Routing [129], NYU Door Opening [130], AUTOLab UR5 [132], and Robot Play [134]. We use the same evaluation and robot embodiment as in the respective publications. For large-scale dataset experiments, we consider Bridge [95] and RT-1 [8] for in-distribution evaluation and use their respective robots: WidowX and Google Robot.
@@ VI Discussion, Future Work, and Open Problems
While RT-X demonstrates a step towards a X-embodied robot generalist, there are many more steps needed to make this future a reality. Our experiments have a number of limitations: it does not consider robots with very different sensing and actuation modalities, it does not study generalization to new robots, and it does not provide a decision criterion for when positive transfer does or does not happen. Studying these questions is an important direction for future work. We hope that this work will serve not only as an example that X-robot learning is feasible and practical, but also provide the tools to advance research in this direction in the future.

### 21 Mar 2024 v5

@@ Abstract
Large, high-capacity models trained on diverse datasets have shown remarkable successes on efficiently tackling downstream applications. In domains from NLP to Computer Vision, this has led to a consolidation of pretrained models, with general pretrained backbones serving as a starting point for many applications. Can such a consolidation happen in robotics? Conventionally, robotic learning methods train a separate model for every application, every robot, and even every environment. Can we instead train “generalist” X-robot policy that can be adapted efficiently to new robots, tasks, and environments? In this paper, we provide datasets in standardized data formats and models to make it possible to explore this possibility in the context of robotic manipulation, alongside experimental results that provide an example of effective X-robot policies. We assemble a dataset from different robots collected through a collaboration between institutions, demonstrating skills ( tasks). We show that a high-capacity model trained on this data, which we call RT-X, exhibits positive transfer and improves the capabilities of multiple robots by leveraging experience from other platforms. The project website is robotics-transformer-x.github.io.
@@ I Introduction
00footnotetext: Allen Institute for AI; Arizona State University; California Institute of Technology; Carnegie Mellon University; Columbia University; EPFL; ETH Zürich; Flexiv Robotics; Georgia Institute of Technology; German Aerospace Center; Google DeepMind; Google Research; IO-AI TECH; Imperial College London; Intrinsic LLC; Istituto Italiano di Tecnologia; Korea Advanced Institute of Science & Technology; Max Planck Institute; Meta AI; Mila Quebec; NVIDIA; New York University; Princeton University; Queensland University of Technology; RIKEN; Shanghai Jiao Tong University; Stanford University; Technische Universität Darmstadt; The University of Texas at Austin; The University of Tokyo; Toyota Research Institute; University of California, Berkeley; University of California, Davis; University of California, San Diego; University of Edinburgh; University of Freiburg; University of Illinois Urbana-Champaign; University of Michigan; University of Montreal; University of Pennsylvania; University of Southern California; University of Technology, Nuremberg; University of Texas at Austin; University of WashingtonA central lesson from advances in machine learning and artificial intelligence is that large-scale learning from diverse datasets can enable capable AI systems by providing for general-purpose pretrained models. In fact, large-scale general-purpose models typically trained on large and diverse datasets can often outperform their narrowly targeted counterparts trained on smaller but more task-specific data. For instance, open-vocab classifiers (e.g., CLIP [1]) trained on large datasets scraped from the web tend to outperform fixed-vocabulary models trained on more limited datasets, and large language models [2, 3] trained on massive text corpora tend to outperform systems that are only trained on narrow task-specific datasets. Increasingly, the most effective way to tackle a given narrow task (e.g., in vision or NLP) is to adapt a general-purpose model. However, these lessons are difficult to apply in robotics: any single robotic domain might be too narrow, and while computer vision and NLP can leverage large datasets sourced from the web, comparably large and broad datasets for robotic interaction are hard to come by. Even the largest data collection efforts still end up with datasets that are a fraction of the size and diversity of benchmark datasets in vision (5-18M) [4, 5] and NLP (1.5B-4.5B) [6, 7]. More importantly, such datasets are often still narrow along some axes of variation, either focusing on a single environment, a single set of objects, or a narrow range of tasks. How can we overcome these challenges in robotics and move the field of robotic learning toward large data regime that has been so successful in other domains?
@@
Following this rationale, we have two goals: (1) Evaluate whether policies trained on data from many different robots and environments enjoy the benefits of positive transfer, attaining better performance than policies trained only on data from each evaluation setup. (2) Organize large robotic datasets to enable future research on X-embodiment models.
@@ V-A In-distribution performance across different embodiments
To assess the ability of RT-X models to learn from X-embodiment data, we evaluate performance on in-distribution tasks. We split our evaluation into two types: evaluation on domains that have small-scale datasets (Fig. 2), where we would expect transfer from larger datasets to significantly improve performance, and evaluation on domains that have large-scale datasets (Table I), where we expect further improvement to be more challenging. Note that we use the same robotics data training mixture (defined in Sec. IV-C) for all the evaluations presented in this section. For small-scale dataset experiments, we use Kitchen Manipulation [128], Cable Routing [129], NYU Door Opening [130], AUTOLab UR5 [132], and Robot Play [134]. We use the same evaluation and robot embodiment as in the respective publications. For large-scale dataset experiments, we consider Bridge [95] and RT-1 [8] for in-distribution evaluation and use their respective robots: WidowX and Google Robot.
@@ VI Discussion, Future Work, and Open Problems
While RT-X demonstrates a step towards a X-embodied robot generalist, many more steps are needed to make this future a reality. Our experiments do not consider robots with very different sensing and actuation modalities. They do not study generalization to new robots, and provide a decision criterion for when positive transfer does or does not happen. Studying these questions is an important future work direction. This work serves not only as an example that X-robot learning is feasible and practical, but also provide the tools to advance research in this direction in the future.